OpenStudy (anonymous):

You roll a pair of number cubes. What is the probability of getting even numbers on both cubes?

OpenStudy (cwrw238):

probability of even number on one cube = 1/2 as the 2 events are independent you multiply the probabilities so your answer is 1/2 * 1/2 = 1/4